OpenStudy (anonymous):

what is the limit of x^4 as x approaches infinity?

OpenStudy (unklerhaukus):

\[\lim\limits_{x\rightarrow\infty}x^4\rightarrow\infty^4=\infty\]

OpenStudy (anonymous):

so if it was asking for a H.A then it would be none... Right?

OpenStudy (unklerhaukus):

i dont know what H.A stands for

OpenStudy (unklerhaukus):

no horizontal asymptotes